A proton has a magnetic field due to its spin on its axis. The field is similar to that created by a circular current loop 0.I got the first part, but am stuck on the 2nd! Thanks in advance!
Answer #1

2) So we know that in rotational kinematics

\tau =I\alpha

here I is moment of inertia of proton which we can treat as solid sphere and

alpha is angular accelaration of proton

and Tau is torque on proton

so,

\tau =\frac{2}{5}mR^2\alpha

so i am assuming torque calculated by you was correct then

\alpha =\frac{5\tau}{2mR^2}

R is radius of proton

m is mass of proton

so plug in all values we get

\alpha =7.21 \times 10^{31}rad/s^2